a dvd player originally cost $150. It is on sale for $72.50. what is the percent of the discount, rounded to the nearest tenth of the percent?
Answer by nerdybill(5404)

a dvd player originally cost $150. It is on sale for $72.50. what is the percent of the discount, rounded to the nearest tenth of the percent?
.
Let x = the discount
then
150 - 150x = 72.50
150(1- x) = 72.50
1-x = 0.48333
-x = -0.5167
x = 0.5167
.
Multiply by 100 to get the percentage:
x = 0.5167*100
x = 51.7%